OpenStudy (anonymous):

What is the solution of log 2x – 3^(125) = 3

OpenStudy (anonymous):

isoloate so that you have log in one side. once you do that then just take log of both sides and solve normally from there.

OpenStudy (anonymous):

not log but an exponent i mean

OpenStudy (anonymous):

What's the base of the log?

OpenStudy (anonymous):

(log base 2x-3) *125 = 3

OpenStudy (anonymous):

\[\large \log_{2x-3} 125 = 3\] \[\large \log_{2x} buwahahaha \]

OpenStudy (anonymous):

forget that second line.. just praticing LaTeXing.. is that first line your problem?

OpenStudy (anonymous):

\[\large \log_{2x-3}{125}=3\]\[\large 125=(2x-3)^3\]\[\large 5=2x-3\]\[\large x=4\]